OpenStudy (anonymous):

Choose the ordered pair that is a solution to the system of equations. 3x - y = 10 -x + y = -4

OpenStudy (anonymous):

Add 3x - y = 10 and -x + y = -4 , you'll get 2x = 6, x= 3.Put the value of x=3 in any equation to get the value of y = -1. So the ordered pair is ( 3 , -1 ).

OpenStudy (anonymous):

What about this one? -3x + y = -4 3x + y = -4

OpenStudy (anonymous):

Add both,you'll get 2y = -8 ,y=-4 and x=0. The ordered pair is (0,-4)